# Correlations for aerodynamic force coefficients of non-spherical particles in compressible flows

### v1

@@ 3.1 Validation
The drag experienced by an isolated spherical particle for particle Reynolds numbers of 100 and 300 is compared with the PR-DNS data of Nagata et al. [29] and the compressible drag model of Loth et al. [24] as shown in Figure 5. For subsonic Mach numbers, results from the current simulations show good agreement with both the PR-DNS data of Nagata et al. [29] and the model of Loth et al. [24]. While the drag coefficient for transonic Mach numbers predicted by our simulations agree better with the data of Nagata et al. [29], the drag coefficient predicted in the supersonic regime is closer to the model of Loth et al. [24]. It is, however, important to note that the transonic regime is not very well documented in the literature and that the existing data shows a lot of variation. Overall, the results for the sphere in compressible flow obtained in this work agree well with the existing literature.
Figure 6 shows a comparison of the simulation results for the prolate spheroid in the incompressible limit ($\mathrm{Ma}=0.3$) for a particle Reynolds number of 300 with the correlations of Sanjeevi et al. [39] and Zastawny et al. [50] for a prolate spheroid in incompressible flow. Furthermore, Figure 7 shows a comparison of the simulation results for the rod-like particle in the incompressible limit ($\mathrm{Ma}=0.3$) for a Reynolds number of 300 with the correlation of Chéron et al. [9] for the same rod-like particle in incompressible flow. The results obtained for the rod-like particle match very well with the correlation of Chéron et al. [9]. The torque coefficient has a slight offset compared to the correlation, which might arise from the fact that different models to calculate the torque coefficient and the centre of pressure have been used. The results obtained for the prolate spheroid do not match the correlations of Zastawny et al. [50] and Sanjeevi et al. [39] perfectly, but are still in reasonable agreement.
@@ 3.2 Non-spherical particles in compressible flow
Figures 8, 9 and 10 show the steady-state velocity flow fields for all three non-spherical particle shapes for a Reynolds number of 300 and Mach numbers of 0.8 and 2, for angles of attack of 0°, 45°, and 90°. These two Mach numbers are representative examples of the flow fields in the transonic ($\mathrm{Ma}=0.8$) and supersonic ($\mathrm{Ma}=2$) regimes. The flow fields of the other Mach numbers do not differ significantly from the two shown. The same applies to the Reynolds number. The flow fields at a Reynolds number of 100 differ for the most part only in magnitude compared to a Reynolds number of 300.
When comparing the velocity fields for $\mathrm{Ma}=2$, as shown in Figures 8, 9 and 10 it becomes clear that the oblate spheroid shape forms the largest bow shock in front of the particle, followed by the rod-like particle, and the prolate spheroid with the smallest bow shock for all three angles. This observation goes hand in hand with the aspect ratios or sphericity of the respective particles. The prolate spheroid shape has the lowest aspect ratio and the fewest “sharp” edges or regions of high curvature, whereas the oblate spheroid shape has the most obvious “sharp” edges.
@@
(a) Prolate spheroid
(b) Oblate spheroid
(c) Rod-like particle
(a) Prolate spheroid
(b) Oblate spheroid
(c) Rod-like particle
Figure 13: Variation of the pitching torque coefficient $C_{\mathrm{M}}$ with Mach number $\mathrm{Ma}$ for different angles of attack considered for the three considered non-spherical particle shapes.
@@
Figures 11, 12 and 13 present the drag coefficients, $C_{\mathrm{D}}$, the lift coefficients, $C_{\mathrm{L}}$, and the pitching torque coefficients, $C_{\mathrm{M}}$, respectively, for all three non-spherical particle shapes as a function of the Mach number for all considered angles of attacks. For angles of attack of 0°and 90°, the lift and torque coefficients are not shown, since they are zero for these flow configurations.
@@
For a given Mach number, the magnitudes of the drag coefficients for all three non-spherical particle shapes in Figure 11 increase with increasing angle of attack. Furthermore, as already indicated by the velocity contour plots, the oblate spheroid with the strongest shock formation has the largest drag coefficients, followed by the rod-like particle, and the prolate spheroid with the weakest shock formation and the smallest drag coefficients. From the literature on airfoils, it is known that an increase in the thickness-to-chord ratio of the airfoil implies an increase in the drag coefficient [1]. A similar behaviour is observed in the case of particles going from a spherical to an elongated shape (from prolate to oblate spheroids). For all three shapes, a smaller particle Reynolds number yields a higher drag coefficient.
@@
For all three non-spherical particle shapes and all angles of attack, the drag coefficient is almost constant or increases only slightly in the subsonic regime as a function of the Mach number, whereas it rapidly increases in the transonic regime and afterwards decreases in the supersonic regime. Intuition suggests that an increase in the Mach number would lead to an increase in the drag coefficient, since the pressure contribution increases for high-speed flows, while the viscous contribution remains significantly smaller. However, in reality, this is only partially the case. The effect of compressibility depends on the Mach number regime of the flow. In the subsonic regime, the drag remains independent of the Mach number up to a critical Mach number, $\mathrm{Ma}_{\mathrm{cr}}$. This critical Mach number is defined as the lowest free-stream Mach number at which the flow around the particle is locally supersonic. If the free-stream Mach number is larger than $\mathrm{Ma}_{\mathrm{cr}}$, the drag starts to increase slowly. A sharp increase in the drag coefficient is observed in the transonic regime followed by a decrease in value for the supersonic regime. For a spherical particle, a similar sharp increase in drag is observed for the transonic regime. However, there is little to no decline of the drag coefficient in the supersonic regimes. The sharp increase of the drag coefficient occurs when the free-stream Mach number is larger than the drag divergence Mach number, which is the Mach number at which the aerodynamic drag begins to increase rapidly as the free-stream Mach number continues to increase. This is due to the formation of a weak shock wave on the surface of the particle, which is preceded by a locally supersonic flow region. This termination of the supersonic flow region by a shock wave gives rise to an adverse pressure gradient. This causes a separation of the boundary layer and a dramatic increase in pressure drag [1]. In the subsonic flow regime, this weak shock wave does not exist, and a plateau region is observed with increasing Mach numbers. The critical Mach number varies for different particle shapes, a blunt body such as a sphere having a relatively high critical Mach number compared to a slender body [1]. Moreover, the extent to which drag decreases after the peak drag is also dependent on the Reynolds number. For a Reynolds number of 100, the drag has a steeper descent than for a Reynolds number of 300.
@@
The lift and pitching torque coefficients in Figures 12 and 13 show a similar Mach number dependence as the previously analysed drag coefficient. The difference in the magnitudes of both force coefficients for different Reynolds numbers is not as large as is the case for the drag coefficient. Overall, the oblate spheroid shows again the largest coefficients for lift and pitching torque, followed by the rod-like particle and prolate spheroid.
@@ 4 Correlations for the aerodynamic force and torque coefficients
The results from the PR-DNS are used to determine the aerodynamic forces and torque, which are used to develop correlations to predict the drag, lift, and torque coefficients for each particle as a function of the particle Reynolds number, $\mathrm{Re_{p}}$, the Mach number, $\mathrm{Ma}$, and the angle of attack between the flow and the major axis of the particle, $\alpha$. The proposed correlations are all valid for a particle Reynolds number that varies in the range $100\leq\mathrm{Re_{p}}\leq 300$, a Mach number of $0\leq\mathrm{Ma}\leq 2$, and an angle of attack of $0\degree\leq\alpha\leq 90\degree$. The correlations to account for compressible effects presented in this section are extensions of the correlations of the coefficients for a particle of the same shape in an incompressible flow, for the same particle Reynolds number and angle of attack. To this end, the correlations for the aerodynamic force coefficients for prolate and oblate spheroids presented by Zastawny et al. [50] and Sanjeevi et al. [39], and the correlations for rod-like particles of Chéron et al. [9] are used as a basis. This assumes that all force coefficients are constant in the range from $\mathrm{Ma}=0$ to $0.3$ for the flow conditions studied. For each correlation, the fitting coefficients for a given function are estimated using curve-fitting algorithms using the Levenberg-Marquardt algorithm of the SciPy library [21]. The accuracy of the new correlations are assessed by calculating the maximum and median errors of their prediction, $\epsilon_{\mathrm{max}}$ and $\epsilon_{\mathrm{med}}$. The error in the prediction is defined as
@@ 4.1 Drag coefficient
$\vartheta_{1}$ $\vartheta_{2}$ $\vartheta_{3}$ $\vartheta_{4}$ $\vartheta_{5}$ $\vartheta_{6}$ $\vartheta_{7}$ $\vartheta_{8}$ $\vartheta_{9}$ $\vartheta_{10}$
Eq. (10) 1.20 $\times$ $10^{-3}$ 0.173 0.721 15.0 0.904 0.171 1.12 0.245 - -
Eq. (11) 1.43 2.00 1.00 $\times$ $10^{-10}$ 4.56 3.87 12.0 0.870 0.159 1.13 0.115
Eq. (12) 1.35 $\times$ $10^{-3}$ 8.54 $\times$ $10^{-2}$ 0.122 1.53 9.49 0.916 8.22 0.709 1.20 0.295
@@ 4.1.1 Prolate spheroid
The PR-DNS results along with the correlation for the prediction of the drag coefficient of the prolate spheroid are shown in the left of Figure 14, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the change in the evolution of the drag coefficient in both transonic and supersonic regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 6.08% and 1.12%, respectively.
@@ 4.1.2 Oblate spheroid
The correlation for the prediction of the drag coefficient for the oblate spheroid is shown in the centre of Figure 14, along with the PR-DNS results. For all angles of attack, except for $\alpha=75\degree$, the correlation accurately recovers the change in the evolution of the drag coefficient in both the transonic and supersonic regimes, for all particle Reynolds numbers. For $\alpha=75\degree$, the variation in the difference between the value of the drag coefficient in the subsonic and supersonic regimes deviates from other angles and remains challenging to capture. However, a very good agreement is observed as the maximum and median prediction errors between the correlation and the PR-DNS are 8.65% and 1.78%, respectively.
@@ 4.1.3 Rod-like particle
$\mathcal{F}_{\mathrm{D}}^{\mathrm{Ma}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}\right)=\frac{\vartheta_{1}\mathrm{Re_{p}}+\vartheta_{2}\left(\mathrm{Re_{p}}\right)^{\vartheta_{3}}\left(\alpha\right)^{\vartheta_{4}}}{\left[1+\exp\left(-\vartheta_{5}\left(\mathrm{Ma}-\vartheta_{6}\right)\right)\right]}+\left({\vartheta_{7}\alpha}\big{/}{\mathrm{Re_{p}}}\right)^{\vartheta_{8}}\exp\left(-\frac{(\mathrm{Ma}-\vartheta_{9})^{2}}{\vartheta_{10}}\right)\,.$ (12)
The first term on the right hand-side of Eq. (12) is similar to the expression given for the prolate spheroid, Eq. (10), as the observations are the same. The larger is the angle of attack, the larger is the difference between the drag coefficient in the subsonic and the supersonic flow regimes. However, the second term on the right hand-side of Eq. (12), which is also based on a Gaussian function, is scaled by the angle of attack. The rationale is that, for the rod-like particle, the bell shape that represents the change in the drag coefficient in the transonic regime is less pronounced for angles of attack close to $\alpha\rightarrow 0\degree$, and is not present for a rod-like particle with an angle of attack $\alpha=0\degree$. This is shown in the right of Figure 14, depicting the prediction of the drag coefficient of the rod-like particle with the correlation along with the PR-DNS results. For all flow cases, a good agreement is observed between the predictions and the PR-DNS results as a maximum and median prediction errors between the correlation and the PR-DNS are 6.93% and 1.44%, respectively.
@@ 4.2 Lift coefficient
Table 2: List of the fitting coefficients in Eqs. (14),  (17), and (18), used in the correlation to predict the change in the drag coefficient in case of compressible flow.
@@ 4.2.1 Prolate spheroid
The PR-DNS results along with the correlation for the prediction of the lift coefficient of the prolate spheroid are shown in the left of Figure 15, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the evolution of the lift coefficient in all regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 10.73% and 0.11%, respectively.

### v2

@@ 3.1 Validation
The drag experienced by an isolated spherical particle for particle Reynolds numbers of 100 and 300 is compared with the PR-DNS data of Nagata et al. [29] and the compressible drag model of Loth et al. [24] as shown in Figure 3. For all Mach numbers, results from the current simulations show good agreement with the PR-DNS data of Nagata et al. [29] and reasonable agreement with the model of Loth et al. [24], while the drag coefficient for supersonic Mach numbers predicted by our simulations agree better with the data of Nagata et al. [29]. It is, however, important to note that the transonic regime is not very well documented in the literature and that the existing data shows a lot of variation. Overall, the results for the sphere in compressible flow obtained in this work agree well with the existing literature.
Figure 4 shows a comparison of the simulation results for the prolate spheroid in the incompressible limit ($\mathrm{Ma}=0.3$) for a particle Reynolds number of 300 with the correlations of Sanjeevi et al. [39] and Zastawny et al. [50] for a prolate spheroid in incompressible flow. Furthermore, Figure 5 shows a comparison of the simulation results for the rod-like particle in the incompressible limit ($\mathrm{Ma}=0.3$) for a Reynolds number of 300 with the data of Chéron et al. [9] for the same rod-like particle in incompressible flow. The results obtained for the rod-like particle match very well with the data of Chéron et al. [9]. The results obtained for the prolate spheroid match the correlations of Zastawny et al. [50] and Sanjeevi et al. [39] very well.
Figure 6 shows the local Mach number and the local Knudsen number for the oblate spheroid at $\mathrm{Re_{p}}=300$ and $\alpha=45\degree$. The oblate spheroid in this flow configuration is one of the most extreme flow conditions in this work. As visible from Figure 6, the local Mach number reaches a maximum of 2.2 and the local particle Knudsen number, computed as [24]
$\mathrm{Kn_{local}}=\sqrt{\frac{\pi\gamma}{2}}\left(\frac{\mathrm{Ma_{local}}}{\mathrm{Re_{local}}}\right)$ (8)
reaches a maximum of 0.096 at the lower tip of the particle surface. Although a part of the flow in the wake of the particle is in the slip regime ($0.10.01<\mathrm{Kn}<$ 0.1), the flow is still dominated by the continuum regime ($\mathrm{Kn}\ll 0.1$). The continuum assupmtion, thus, forms an adequate basis for the conducted simulations.
Figure 6: Contour plots of the local Mach number (left) and of the local particle Knudsen number (right) for the oblate spheroid at $\mathrm{Re_{p}}=300$ and $\alpha=45\degree$.
@@ 3.2 Non-spherical particles in compressible flow
Figures 7, 8 and 9 show the steady-state velocity flow fields for all three non-spherical particle shapes for a Reynolds number of 300 and Mach numbers of 0.8 and 2, for angles of attack of 0°, 45°, and 90°. These two Mach numbers are representative examples of the flow fields in the transonic ($\mathrm{Ma}=0.8$) and supersonic ($\mathrm{Ma}=2$) regimes. The flow fields of the other Mach numbers do not differ significantly from the two shown. The same applies to the Reynolds number. The flow fields at a Reynolds number of 100 differ for the most part only in magnitude compared to a Reynolds number of 300.
When comparing the velocity fields for $\mathrm{Ma}=2$, as shown in Figures 7, 8 and 9 it becomes clear that the oblate spheroid shape forms the largest bow shock in front of the particle, followed by the rod-like particle, and the prolate spheroid with the smallest bow shock for all three angles. This observation goes hand in hand with the aspect ratios or sphericity of the respective particles. The prolate spheroid shape has the lowest aspect ratio and the fewest “sharp” edges or regions of high curvature, whereas the oblate spheroid shape has the most obvious “sharp” edges.
@@
Figure 10 presents the drag coefficients, $C_{\mathrm{D}}$, for all three non-spherical particle shapes as a function of the Mach number for all considered angles of attack. The lift and torque coefficients as a function of the Mach number for all three non-spherical particle shapes are shown in B.
@@
For a given Mach number, the magnitudes of the drag coefficients for all three non-spherical particle shapes in Figure 10 increase with increasing angle of attack. Furthermore, as already indicated by the velocity contour plots, the oblate spheroid with the strongest shock formation has the largest drag coefficients, followed by the rod-like particle, and the prolate spheroid with the weakest shock formation and the smallest drag coefficients. For all three shapes, a smaller particle Reynolds number yields a higher drag coefficient.
@@
For all three non-spherical particle shapes, the drag coefficient in the subsonic regime increases only slightly for angles of attack of 0°to 15°, but increases more significantly as the angle of attack increases, as a function of the Mach number. For all angles of attack, the drag coefficient increases even more rapidly in the transonic regime and, afterwards, decreases in the supersonic regime. Intuition suggests that an increase in the Mach number leads to an increase in the drag coefficient, since the pressure contribution increases for high-speed flows, while the viscous contribution remains significantly smaller. However, in reality, this is only partially the case. The effect of compressibility depends on the Mach number regime of the flow. In the subsonic regime, the drag coefficient remains independent of the Mach number up to a critical Mach number, $\mathrm{Ma}_{\mathrm{cr}}$. This critical Mach number is defined as the lowest free-stream Mach number at which the flow around the particle is locally supersonic. If the free-stream Mach number is larger than $\mathrm{Ma}_{\mathrm{cr}}$, the drag coefficient starts to increase slowly. A sharp increase in the drag coefficient is observed in the transonic regime followed by a decrease in value for the supersonic regime. For a spherical particle, a similar sharp increase in the drag coefficient is observed for the transonic regime. However, there is little to no decline of the drag coefficient in the supersonic regimes. The sharp increase of the drag coefficient occurs when the free-stream Mach number exceeds a critical value known as the drag divergence Mach number, which marks the onset of a rapid increase in aerodynamic drag due to the formation of shock waves and compressibility effects. This termination of the supersonic flow region by a shock wave gives rise to an adverse pressure gradient. This causes a separation of the boundary layer and a dramatic increase in pressure drag [1]. The critical Mach number varies for different particle shapes, a blunt body such as a sphere having a relatively high critical Mach number compared to a slender body [1]. Moreover, the extent to which the drag coefficient decreases after the peak drag is also dependent on the Reynolds number. For a Reynolds number of 100, the drag coefficient has a steeper descent than for a Reynolds number of 300.
@@
The lift and torque coefficients shown in Figures 18 and 19 show a similar Mach number dependence as the previously analysed drag coefficient. The difference in the magnitudes of both force coefficients for different Reynolds numbers is not as large as is the case for the drag coefficient. Overall, the oblate spheroid shows the largest coefficients for lift and torque, followed by the rod-like particle and prolate spheroid. The torque coefficient has, in comparison to the drag and lift coefficients, a much steeper descent in the supersonic regime and the torque coefficients in the supersonic flow regime are even smaller compared to the coefficients in the subsonic regime.
@@ 4 Correlations for the aerodynamic force and torque coefficients
The results from the PR-DNS are used to determine the aerodynamic forces and torque, which are used to develop correlations to predict the drag, lift, and torque coefficients for each particle as a function of the particle Reynolds number, $\mathrm{Re_{p}}$, the Mach number, $\mathrm{Ma}$, and the angle of attack between the flow and the major axis of the particle, $\alpha$. The proposed correlations are all valid for a particle Reynolds number that varies in the range $100\leq\mathrm{Re_{p}}\leq 300$, a Mach number of $0\leq\mathrm{Ma}\leq 2$, and an angle of attack of $0\degree\leq\alpha\leq 90\degree$. The correlations to account for compressible effects presented in this section are extensions of the correlations of the coefficients for a particle of the same shape in an incompressible flow, for the same particle Reynolds number and angle of attack ($C_{\mathrm{D,L,M}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}=0\right)$). To this end, for instance, the correlations for the aerodynamic force coefficients for prolate and oblate spheroids presented by Zastawny et al. [50] and Sanjeevi et al. [39], and the correlations for rod-like particles of Chéron et al. [9] and Zastawny et al. [50] could be used as a basis for $C_{\mathrm{D,L,M}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}=0\right)$. This assumes that all force coefficients are constant in the range from $\mathrm{Ma}=0$ to $0.3$ for the flow conditions studied. In C a summary of the full models for all three non-spherical particles shapes are given exemplary with the correlations of Zastawny et al. [50] for the incompressible limit. For each correlation, the fitting coefficients for a given function are estimated using curve-fitting algorithms using the Levenberg-Marquardt algorithm of the SciPy library [21]. The accuracy of the new correlations are assessed by calculating the maximum and median errors of their prediction, $\epsilon_{\mathrm{max}}$ and $\epsilon_{\mathrm{med}}$. The error in the prediction is defined as
@@ 4.1 Drag coefficient
$\vartheta_{1}$ $\vartheta_{2}$ $\vartheta_{3}$ $\vartheta_{4}$ $\vartheta_{5}$ $\vartheta_{6}$ $\vartheta_{7}$ $\vartheta_{8}$ $\vartheta_{9}$ $\vartheta_{10}$ $\vartheta_{11}$
Eq. (11) 1.55 $\times$ $10^{-3}$ 0.201 0.492 15.6 0.878 0.183 1.21 0.142 - - -
Eq. (12) 6.07 0.269 4.40 $\times$ $10^{-5}$ 2.10 3.90 21.3 0.853 0.130 1.17 0.252 -
Eq. (13) 1.14 $\times$ $10^{-3}$ 0.121 2.64 0.0983 1.40 14.6 0.861 3.94 0.631 1.20 0.330
@@ 4.1.1 Prolate spheroid
The PR-DNS results along with the correlation for the prediction of the drag coefficient of the prolate spheroid are shown in the left of Figure 11, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the change in the evolution of the drag coefficient in both transonic and supersonic regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 7.56% and 1.77%, respectively.
@@ 4.1.2 Oblate spheroid
The correlation for the prediction of the drag coefficient for the oblate spheroid is shown in the centre of Figure 11, along with the PR-DNS results. For all angles of attack, except for $\alpha=75\degree$, the correlation accurately recovers the change in the evolution of the drag coefficient in both the transonic and supersonic regimes, for all particle Reynolds numbers. For $\alpha=75\degree$, the variation in the difference between the value of the drag coefficient in the subsonic and supersonic regimes deviates from other angles and remains challenging to capture. However, a very good agreement is observed as the maximum and median prediction errors between the correlation and the PR-DNS are 10.19% and 2.1%, respectively.
@@ 4.1.3 Rod-like particle
${{\mathcal{F}_{\mathrm{D}}^{\mathrm{Ma}}\left(\mathrm{Re_{p}},\alpha,\mathrm{Ma}\right)=\frac{\vartheta_{1}\cdot\mathrm{Re_{p}}+\vartheta_{2}\cdot\left(\frac{\mathrm{Re_{p}}}{\vartheta_{3}}\right)^{\vartheta_{4}}\cdot\alpha^{\vartheta_{5}}}{\left[1.0+\exp{\left(-\vartheta_{6}\cdot(\mathrm{Ma}-\vartheta_{7})\right)}\right]}+\left(\frac{\vartheta_{8}\cdot\alpha}{\mathrm{Re_{p}}}\right)^{\vartheta_{9}}\cdot\exp{\left(-\frac{(\mathrm{Ma}-\vartheta_{10})^{2}}{\vartheta_{11}}\right)}}}$ (13)
The first term on the right hand-side of Eq. (13) is similar to the expression given for the prolate spheroid, Eq. (11), as the observations are the same. The larger is the angle of attack, the larger is the difference between the drag coefficient in the subsonic and the supersonic flow regimes. However, the second term on the right hand-side of Eq. (13), which is also based on a Gaussian function, is scaled by the angle of attack. The rationale is that, for the rod-like particle, the bell shape that represents the change in the drag coefficient in the transonic regime is less pronounced for angles of attack close to $\alpha\rightarrow 0\degree$, and is not present for a rod-like particle with an angle of attack $\alpha=0\degree$. This is shown in the right of Figure 11, depicting the prediction of the drag coefficient of the rod-like particle with the correlation along with the PR-DNS results. For all flow cases, a good agreement is observed between the predictions and the PR-DNS results as a maximum and median prediction errors between the correlation and the PR-DNS are 7.06% and 1.29%, respectively.
@@ 4.2 Lift coefficient
Table 2: List of the fitting coefficients in Eqs. (15),  (18), and (19), used in the correlation to predict the change in the lift coefficient in case of compressible flow.
@@ 4.2.1 Prolate spheroid
The PR-DNS results along with the correlation for the prediction of the lift coefficient of the prolate spheroid are shown in the left of Figure 12, as a function of the Mach number, for all angles of attack and particle Reynolds number values considered in this work. The correlation accurately recovers the evolution of the lift coefficient in all regimes, for all particle Reynolds numbers and angles of attack. The maximum and median prediction errors between the correlation and the PR-DNS are 13.93% and 0.72%, respectively.